OpenStudy (anonymous):

what the common difference between these fractions. 2/5, 9/10, 7/5, 19/10, 12/5

OpenStudy (anonymous):

Subtract the first term from the second term.\[\frac{9}{10} - \frac25\]\[\frac{9}{10} - \frac{4}{10} = \frac{5}{10} = \frac12\]

OpenStudy (anonymous):

oohh so you have to change the denominator??